# Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles

### v2

@@ 3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning-preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix E, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 5 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix A, Table 11.
@@ 3.4 Data Quality Validation
Table 5: Descriptive statistics for SciTrek (test set). Length: the length level of input tokens; Instances: number of question-answer pairs with their contexts; Articles/C: average number of scientific articles per context; Words/Q: average number of words per question; Words/A: average number of words per answer.
@@
Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 7, Appendix E). In each annotation session, crowdworkers were given three questions with different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session and were compensated above the UK living wage, at £12 per hour.
@@
Table 7: Evaluation of open-weight (top) and proprietary (bottom) models on SciTrek using exact match (%). Results are reported for two settings: using the full-text scientific articles as context and using the corresponding database tables. F-* refers to test data showing full-text article clusters with different maximum lengths, 64K/128K/512K/1M. D-* refers to test data showing the database tables corresponding to F-64/128/512/1024 as input. Note that for any given article cluster, the database representation is significantly shorter (mean length $\sim$}2K tokens); therefore we are able to pass database representations for all lengths to all models. Context Size denotes the maximum context length supported by each model. — indicates the model cannot handle the given context size; $\star$ indicates that the model was not evaluated due to prohibitive computational cost.
4 Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For each family, we select the largest model we can feasibly run with our resources. [^13] Experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1. Table 7 provides an overview of the models we consider, which vary in parameter scale and supported context length (most were released within the past six months). For detailed model descriptions and settings, refer to Appendix B. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles. For higher confidence, reported metrics are averages across three runs.
@@ 4.1 Zero-shot Prompting
Our zero-shot results are summarized in Table 7. Overall, we observe that SciTrek is challenging especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones. We observe similar tendencies when using F1 as the evaluation metric (see Appendix C).
@@
Figure 2: Fine-grained analysis for best performing models in terms of exact match (EM). We examine how model performance varies across question topics (top), information processing skills (middle) and subjects (bottom) when using full-text scientific articles as context with an input length of 128K.
@@ 4.2 Long-context Post-training
Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths including 64K, 128K, 512K, and 1M. We use this data to assess whether more training could improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M on data with maximum context of 128K tokens (7,703 training instances).
Moreover, we examine whether the post-trained model can generalize along the dimensions of input length, question topics, and information processing skills. We compare in-distribution performance where the model is trained and tested in similar conditions (e.g., on lengths of 64K and 128K) to out-of distribution where the model is tested on an unseen dimension. Specifically, we train the model on data corresponding to 64K and 128K context lengths and evaluate it both in-distribution (64K, 128K) and out-of-distribution (512K, 1M). For question topic, training uses author- and title-related questions at 64K and 128K, with evaluation covering the same topics in-distribution and citation-related questions out-of-distribution. For skills, training is performed on non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, Filtering+Aggregating), with evaluation including these in-distribution skills and the out-of-distribution Relational Filtering skill.
For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of $2\times 10^{-6}$, and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung and Lapata, 2025; Zheng et al., 2025). We found a simple sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. The prompt for generating reasoning traces is provided in Appendix E. Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
@@
Table 8: Post-training Qwen2.5-7B-Instruct-1M on SciTrek. Results with SFT and GRPO are compared with zero-shot prompting (ZS) of the same model and other proprietary models in terms of exact match (%). ID is a shorthand for in-distribution, and OOD for out-of-distribution. —: the model cannot handle the context size or has prohibitive computational cost.
The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size. However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on out-of-distribution question topics and information processing skills.
5 Model Failure Analysis
Model #Articles Question Length SQL Length Ground-truth Answer Length
@@ 5 Model Failure Analysis
Table 9: Pearson correlation between various factors and model performance using exact match (with full-text articles as context, 128K token length). The length of questions, SQL queries, and ground-truth answers is computed by counting words separated by spaces. Bold values indicate the strongest correlation for each model, and denotes significant correlations with p-value <<0.05. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block.
We conduct additional analysis to understand model failures on SciTrek when using full-text articles as context. Table 9 shows correlations between model performance and various factors: the number of input articles, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA). They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the number of input articles and the length of the SQL query.
Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1). More details of failure analysis are provided in Appendix D.
@@
We also manually analyzed 200 reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. [^14] Reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B are lengthy, averaging 638, 1,685, and 2,320 respectively. We will release these traces along with our code, but we omit them here due to space constraints. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. Specifically, we examined whether the reasoning is logically sound and whether individual steps obtain accurate information from specific contexts. We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths for all three models, although their traces occasionally include repeated steps or extraneous operations unrelated to the question; (2) despite not being trained for 512K, Qwen2.5-7B-1M (GRPO) still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers from these models especially for Qwen2.5-7B-1M (GRPO). These findings indicate that GRPO-based reinforcement learning improves abstract reasoning without enhancing the model’s accuracy in fine-grained operations, such as counting references.

### v3

@@ 3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning-preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix G, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 5 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix B, Table 17.
@@ 3.4 Data Quality Validation
Table 5: Descriptive statistics for the SciTrek test set. Length: the length level of input tokens; Instances: number of question-answer pairs with their contexts; Articles/C: average number of articles per context; Words/Q: average number of words per question; Words/A: average number of words per answer.
@@
Each instance was annotated by three participants following the same instructions as those used for model testing with database tables as context (Figure 7, Appendix G). In each annotation session, crowdworkers were presented with three questions over different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session and were compensated above the UK living wage, at £12 per hour.
@@
Table 7: Evaluation of SQL execution-based method (top), open-weight (middle) and proprietary (bottom) models on SciTrek using exact match (%). Results are reported for two context settings: full-text scientific articles and corresponding database tables. F-* denotes test instances using full-text article clusters of up to 64K, 128K, 512K, and 1M tokens. D-* denotes test data instances using the database tables corresponding to F-64, 128, 512, and 1024. For any given article collection, the database representation is substantially shorter (mean $\sim$}2K tokens), and can therefore be passed to all models regardless of their context limit. Context Size denotes the maximum input length supported by each model. — indicates the model cannot process the given context size; $\star$ indicates the model was not evaluated due to prohibitive computational cost.
4 Experimental Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For proprietary models, we include o4-mini [^10] https://developers.openai.com/api/docs/models/o4-mini and Gemini 2.5 Pro (Gemini Team, 2025), which are widely regarded as the leading models for long-context reasoning. For open-weight models, we select the largest model we can feasibly run with our resources from each model family. [^11] Experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct (Llama Team, 2025), and the distilled variant of DeepSeek-R1 (DeepSeek-AI et al., 2024). Table 7 provides an overview of the models we consider, which vary in parameter scale and supported context length (most were released within the past six months). For detailed model descriptions and settings, refer to Appendix C. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context, and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix G. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles. For higher confidence, reported metrics are averages across three runs.
@@ 4.1 Zero-shot Prompting
We evaluate zero-shot prompting performance on our test data; results are summarized in Table 7. Overall, we observe that SciTrek is challenging to most experimental models, especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones.
We also evaluate SQL execution-based methods (see the first row in Table 7). We first use Qwen2.5-Coder-32B-Instruct to convert the natural language questions into SQL queries, and then execute these on the corresponding database tables. When the input consists of full-text articles, we construct database tables based on the retrieved titles, authors, and references using the same model. [^12] The prompts for extracting titles, authors, and references are in Appendix D. The results show that retrieval-augmented SQL execution performs poorly, while SQL execution on gold database tables provides an upper bound performance of all retrieval-augmented methods. We observe similar tendencies when using F1 as the evaluation metric (see Appendix E).
@@
Figure 2: Fine-grained analysis for best performing models in terms of exact match (EM). We examine how model performance varies across question topics (top), information processing skills (middle) and subjects (bottom) when using full-text scientific articles as context at an input length of 128K tokens.
@@ 4.2 Long-context Post-training
Using the data construction methodology described in Section 3, we curate a substantial training dataset (19,543 instances) spanning four context lengths: including 64K, 128K, 512K, and 1M. We use this data to investigate whether additional training can improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M trained on contexts of up to 128K tokens (7,703 training instances).
We further examine whether the post-trained model can generalize along the dimensions of input length, question topic, and information processing skills. For each dimension, we compare in-distribution performance, where training and testing conditions are matched, against out-of-distribution performance where the model is evaluated on an unseen dimension. For input length, the model is trained on contexts not exceeding 128K tokens and evaluated both within this range (in-distribution) and on longer contexts (out-of-distribution). For question topic, training uses author- and title-related questions at 64K and 128K tokens, with in-distribution evaluation on the same topics and out-of-distribution evaluation on citation-related questions. For skills, training covers non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, and Filtering+Aggregating), with out-of-distribution evaluation on the held-out Relational Filtering skill.
For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of $2\times 10^{-6}$, and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung and Lapata, 2025; Zheng et al., 2025). We found a simple sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. [^13] The prompt for generating reasoning traces is provided in Appendix G. Since GRPO optimization is time-consuming, we restrict training to a single epoch, which takes around 5 days.
@@
Table 8: Post-training results for Qwen2.5-7B-Instruct-1M on SciTrek. We compare SFT and GRPO against zero-shot prompting (ZS) of the same model and selected proprietary models, reporting exact match (%). ID is a shorthand for in-distribution, and OOD for out-of-distribution. —: the model cannot process the given context size, or evaluation was not feasible due to prohibitive computational cost.
Models LongBench v2 Loong
Qwen2.5 (ZS) 34.3 33.3
Qwen2.5 (SFT) 35.5 39.3
Qwen2.5 (GRPO) 38.4 34.7
Table 9: Evaluation results on LongBench v2 and Loong. Performance is reported as single-choice question accuracy (%) and answer quality (scored on a 1–100 scale by GPT-5 mini as judge), respectively; higher values indicate better performance. We follow the evaluation protocols of  Bai et al. (2025) and Wang et al. (2024c).
Model #Articles Question Length SQL Length Gold Answer Length
@@
Table 10: Pearson correlation between various factors and model performance using exact match (with full-text articles as context, 128K token length). The length of questions, SQL queries, and ground-truth answers is computed by counting words separated by spaces. Bold values indicate the strongest correlation per model, and denotes significant correlations with p-value <<0.05. The first block shows zero-shot models; the second block shows supervised variants of Qwen2.5-7B-Instruct-1M.
The exact match results in Table 8 show that both SFT and GRPO yield modest improvements across all dimensions, with the post-trained model surpassing GPT-4.1 despite its considerably larger parameter count. Overall performance nonetheless remains weak: the best absolute score (22.5%) falls well below both Gemini (28.1%) and o4-mini (48.4%). Furthermore, while Qwen2.5-7B-Instruct-1M still fails to generalise to longer inputs, post-training does improve out-of-distribution performance on question topics and information-processing skills.
We also test the post-trained models on existing out-of-domain benchmarks, including LongBench v2 and Loong, where numerical reasoning is implicitly required. The results in Table 9 show that post-training on SciTrek can also improve model performance on these broader long-context benchmarks, suggesting that the acquired skills transfer beyond the training domain.
Model A S F F+A F+S R
Qwen2.5-7B-Instruct-1M 90 10 80 10 40 70
Qwen2.5-14B-Instruct-1M 0 0 30 30 30 80
Qwen3-4B-Instruct-2507 50 60 80 10 90 90
Qwen3-4B-Thinking-2507 0 0 20 20 10 40
Qwen3-30B-A3B-Inst.-2507 0 0 0 20 0 70
Qwen3-30B-A3B-Thin.-2507 0 0 0 30 10 30
Gemma-3-27B-IT 10 0 0 40 10 40
Llama-4-Scout-Instruct 10 10 10 60 30 80
Llama-3.3-70B-Instruct 20 30 40 70 30 70
DeepSeek-R1-Dist.-Llama-70B 0 0 10 10 0 0
Gemini 2.5 Pro 0 0 0 20 0 30
GPT-4.1 0 0 0 0 0 10
o4-mini 0 0 0 30 10 20
Qwen2.5-7B-Inst.-1M (SFT) 0 0 20 70 20 0
Qwen2.5-7B-Inst.-1M (GRPO) 0 0 0 0 0 0
Table 11: Proportion of samples (%) where models output output “NULL”, broken down by information processing skill. The first block shows zero-shot models; the second block shows supervised variants of Qwen2.5-7B-Instruct-1M. A abbregivats Aggregation, S stands for Sorting, F for Filtering, and R for Relational Filtering.
Model A S F F+A F+S R
Qwen2.5-7B-Instruct-1M 0 0 0 0 0 10
Qwen2.5-14B-Instruct-1M 10 30 0 20 30 10
Qwen3-4B-Instruct-2507 0 0 0 0 0 0
Qwen3-4B-Thinking-2507 0 0 0 0 0 0
Qwen3-30B-A3B-Inst.-2507 30 0 50 40 10 10
Qwen3-30B-A3B-Thin.-2507 0 10 0 0 0 0
Gemma-3-27B-IT 20 30 40 10 0 30
Llama-4-Scout-Instruct 10 30 10 20 0 0
Llama-3.3-70B-Instruct 0 10 10 10 10 10
DeepSeek-R1-Dist.-Llama-70B 0 0 0 0 0 0
Gemini 2.5 Pro 10 0 0 10 0 0
GPT-4.1 50 40 40 30 10 10
o4-mini 10 10 0 0 0 0
Qwen2.5-7B-Inst.-1M (SFT) 0 0 0 0 0 0
Qwen2.5-7B-Inst.-1M (GRPO) 10 0 0 0 10 0
Table 12: Proportion of samples (%) where models do not follow the specified answer format, broken down by information processing skill. The first block shows zero-shot models; the second block shows supervised versions of Qwen2.5-7B-Instruct-1M. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and R for Relational Filtering.
Model A S F F+A F+S R
Qwen2.5-7B-Instruct-1M 0 0 10 0 0 0
Qwen2.5-14B-Instruct-1M 0 0 0 0 0 10
Qwen3-4B-Instruct-2507 0 0 0 0 0 0
Qwen3-4B-Thinking-2507 0 0 0 10 0 10
Qwen3-30B-A3B-Inst.-2507 0 0 0 0 10 0
Qwen3-30B-A3B-Thin.-2507 10 0 30 30 10 30
Gemma-3-27B-IT 10 10 0 0 0 0
Llama-4-Scout-Instruct 0 0 0 0 0 0
Llama-3.3-70B-Instruct 10 0 20 0 10 0
DeepSeek-R1-Dist.-Llama-70B 0 0 0 10 10 10
Gemini 2.5 Pro 20 0 10 20 0 10
GPT-4.1 0 0 10 0 10 10
o4-mini 20 10 30 10 0 0
Qwen2.5-7B-Inst.-1M (SFT) 0 0 10 0 0 20
Qwen2.5-7B-Inst.-1M (GRPO) 10 20 10 10 0 10
Table 13: Proportion of samples (%) where models generate partial answers broken down by information processing skill. The first block shows zero-shot models; the second lock shows supervised versions of Qwen2.5-7B-Instruct-1M. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and R for Relational Filtering.
Model F F+A F+S R All
Qwen2.5-7B-Instruct-1M 0 0 0 0 0
Qwen2.5-14B-Instruct-1M 0 20.0 0 0.2 3.8
Qwen3-4B-Instruct-2507 0 0 0 0 0
Qwen3-4B-Thinking-2507 0 53.3 8.3 1.1 9.1
Qwen3-30B-A3B-Inst.-2507 0 0 0 0 0
Qwen3-30B-A3B-Thin.-2507 0 66.7 16.7 1.1 12.9
Gemma-3-27B-IT 0 6.7 0 1.1 1.5
Llama-4-Scout-Instruct 0 13.3 0 2.1 3.0
Llama-3.3-70B-Instruct 0 6.7 0 0 0.8
DeepSeek-R1-Dist.-Llama-70B 0 13.3 0 0 3.0
Gemini 2.5 Pro 0 20.0 8.3 6.4 9.8
GPT-4.1 0 13.3 0 3.2 4.5
o4-mini 0 60.0 50.0 5.3 18.9
Qwen2.5-7B-Inst.-1M (SFT) 0 6.7 0 10.6 9.1
Qwen2.5-7B-Inst.-1M (GRPO) 0 33.3 8.3 4.3 8.3
Table 14: Model performance on questions involving negation (e.g., “not" or “never"), reported in terms of exact match (%). The first block shows zero-shot models; the second block shows supervised versions of Qwen2.5-7B-Instruct-1M. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and R for Relational Filtering.
5 Model Failure Analysis
What factors correlate with model failures?
Table 10 reports correlations between model performance and four factors: the number of input articles (#Articles), the question length in words, the length of the underlying SQL query in words, and the length of the ground-truth answer in words. Across most models, performance correlates primarily with question length, whereas o4-mini additionally exhibits correlations with the number of input articles and the SQL query length.
What failure patterns emerge from model outputs?
We randomly selected 60 instances at 128K input length where all zero-shot and supervised models fail (10 instances per skill in Table 4) and identify four recurring failure patterns.
Fallback to NULL. As shown in Table 11, many open-weight models frequently output “NULL”, particularly on filtering tasks. This suggests they rely on the fallback instruction [^14] Our instruction prompts models to generate “NULL” when they could not find an answer in the context. rather than genuinely reasoning over the provided context.
Format non-compliance. Table 12 shows that models often fail to follow the specified answer format, especially for sorting-related questions, for instance, generating author lists when sorted author counts are requested, or returning lists when scalar aggregates are expected. Notably, GPT-4.1 exhibits format violations far more frequently than any other model.
Partial answers. Table 13 reveals that some models produce incomplete answers when lists are requested, particularly on aggregation tasks. Qwen3-30B-A3B-Thinking-2507 is the most prone to this behaviour among all evaluated models.
Negation-based filtering. Table 14 shows that all models struggle with questions involving negation (e.g., those containing “not” or “never”). Our test set includes 132 such instances across filtering-related skills (Filtering: 11, Filtering+Aggregation: 15, Filtering+Sorting: 12, Relational Filtering: 94), and performance is consistently poor across the board.
What do reasoning traces reveal about failure modes?
@@ 5 Model Failure Analysis
We manually analysed 200 reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B across 25 randomly sampled SQL query templates. [^15] Reasoning traces from these three models are lengthy, averaging 638, 1,685, and 2,320 tokens respectively. We release the traces alongside our code; they are omitted here due to space constraints. Reasoning traces from Gemini 2.5 Pro, GPT-4.1, and o4-mini are not accessible for analysis. For each trace, we assessed whether the reasoning is logically coherent and whether individual steps retrieve accurate information from the context. Three findings stand out. First, reasoning remains largely coherent at 64K and 128K context lengths across all three models, though traces occasionally include repeated or extraneous steps. Second, despite not being trained on 512K contexts, Qwen2.5-7B-1M (GRPO) still produces coherent reasoning structure at that length, albeit with reduced step-level accuracy compared to shorter contexts. Third, steps requiring counting—particularly of references—are frequently inaccurate, leading to incorrect final answers, most markedly for Qwen2.5-7B-1M (GRPO). Taken together, these findings suggest that GRPO-based reinforcement learning improves abstract reasoning while leaving fine-grained numerical operations, such as counting references, largely unaddressed. Further details are provided in Appendix F.